suppose that 500 people attended the play if adults were charged 10 dollars and children were charged 7 dollars determine how many adults and children attended if they raised a total of 4472 dollars
Answer by nerdybill(7384)

suppose that 500 people attended the play if adults were charged 10 dollars and children were charged 7 dollars determine how many adults and children attended if they raised a total of 4472 dollars
.
Let x = number of adults
then
500-x = number of children
.
10x + 7(500-x) = 4472
10x + 3500-7x = 4472
3x + 3500 = 4472
3x = 972
x = 324 (adults)
.
Children:
500-x = 500-324 = 176